OpenStudy (anonymous):

HELP PLEASE!! MEDALS!! Atoms contain charged particles called protons and electrons. Each proton has a charge of +1, and each electron has a charge of –1. A sulfur ion has 16 protons and 18 electrons. Find the net charge.

OpenStudy (nubeer):

hmm this is a chemistry question not math one. try posting it in there u might get answer.

OpenStudy (anonymous):

no it's actually algebra

OpenStudy (nubeer):

ahh it's really not.

OpenStudy (nubeer):

net charge = no. of protons(charge of protons) + no of electrons *( charge of electrons.)

OpenStudy (anonymous):

is it 34

OpenStudy (anonymous):

17+17

OpenStudy (anonymous):

16 +1 =17 18-1= 17 34 the net charge

OpenStudy (nubeer):

nooo.. how u even got 17?

OpenStudy (nubeer):

16(+1) + 18(-1)

OpenStudy (anonymous):

oh each proton has a charge of +1

OpenStudy (anonymous):

so why would this question be in my pre-Algebra book?

OpenStudy (anonymous):

+2?

OpenStudy (anonymous):

i believe is 34

OpenStudy (nubeer):

lol no clue.. i use to do these one in chemistry.. and no not +2 it's -2

OpenStudy (agent0smith):

It's an algebra question on adding positive and negative integers.

OpenStudy (anonymous):

it's*

OpenStudy (anonymous):

numbeer you are busting my chops around this question!..lol

OpenStudy (agent0smith):

16(1) + 18(-1) =

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

it that right?

OpenStudy (anonymous):

@Best_Mathematician

OpenStudy (anonymous):

@KeganMoore

OpenStudy (anonymous):

@Hero

OpenStudy (nubeer):

it is -2

OpenStudy (anonymous):

@Hero can you please confirm my answer

OpenStudy (agent0smith):

@nubeer already said the answer earlier: "nubeer Best Response Undo 2 lol no clue.. i use to do these one in chemistry.. and no not +2 it's -2" And then confirmed it again "nubeer Best Response Undo 2 it is -2" Yes, it's -2. 16-18=-2